# mOSCAR: A Large-scale Multilingual and Multimodal Document-level Corpus

### v1

@@ 5.2 Results
Table 5: Translate-test results averaged over languages.
Figure 4: Score differences averaged over benchmarks and languages between the model trained on mOSCAR + text-image pairs and the model trained only on text-image pairs.
@@
6 Conclusion, Limitations and Societal Impacts
We introduce mOSCAR, a large-scale multilingual and multimodal dataset covering 163 languages and composed of 315M documents, 214B tokens and 1.2B images. We show that mOSCAR is of good quality, diverse and can be used to train a multilingual and multimodal LLM. We ensure that mOSCAR is as safe as possible by applying a series of filtering steps to remove NSFW content. We however did not conduct any toxicity analysis or evaluate its biases as this is challenging in a multilingual setting. As it is crawled from the internet, it is possible that mOSCAR reflects biases widespread on it. Nevertheless, by its multilingual nature, mOSCAR is a step towards the inclusion of more languages, cultures and people in accessing mLLMs.
@@ Acknowledgements
This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011014232R1, 2023-AD011014232 and 2023-AD011012254 made by GENCI. It was also partly funded by the last three authors’ chairs in the PRAIRIE institute funded by the French national agency ANR as part of the “Investissements d’avenir” programme under the reference ANR-19- P3IA-0001. We deeply thanks the Jean-Zay support team. We also thanks Filip Šedivý for insightful discussions regarding the removal of CSAM, Thorn for having provided access to their CSAM detector, Zeeshan Khan for discussions regarding the training of the models and Victoria Le Fourner for having manually checked subsamples of NSFW images.

### v2

@@ 5.1 Evaluation setup
Table 8: Translate-test results averaged over languages where all benchmarks were translated from local languages into English using Google Translate API. Multi. OF mOSCAR + cap. refers to Multilingual Open Flamingo trained on mOSCAR and text-image pairs while OF-3B MPT refers to Open Flamingo (Awadalla et al., 2023) based on MPT (Team, 2023) and trained on mmc4 (Zhu et al., 2023) and English text-image pairs.
Figure 4: Score differences averaged over benchmarks and languages between the model trained on mOSCAR + text-image pairs and the model trained only on text-image pairs. Bold is best result.
5.2 Results
Tables 5 and 8 show the average results across all languages. Full results are available in Appendix A.7. We notice that the multilingual OpenFlamingo trained additionally on mOSCAR gets better results than the model trained on captioning data only while having seen fewer image-text pairs during training. More importantly, when increasing the number of few-shot examples from 0 to 16, it sees gains of on average +6.71 points on VQA benchmarks and +19.39 CideR points on captioning benchmarks. In contrast, the model trained on text-image pairs only sees gains of +2.82 and +9.08 points respectively. In cross-modal machine translation, the model additionally trained on interleaved data is again much better than the one trained on just captioning data, which is not able to translate the Multi30k benchmark at all. [^20] Most of the time, the model is not able to follow the prompt and only outputs the end of sequence token. Moreover, mOSCAR helps the model to learn to zero-shot disambiguate translations as shown by the improved average score on CoMMuTE (63.54) compared to the model trained on captions only (61.36).
Multilingual Open Flamingo trained on mOSCAR & text-image pairs is also better than OpenFlamingo 3B MPT evaluated on translate test benchmarks [^21] This means benchmarks were translated from local languages to English, using Google Translate API. However, we obtain the best results (except for MaXM) by evaluating our multilingual Open Flamingo on the translate-test benchmarks since the underlying language model (Gemma-2B) is far better in English than other languages. We also notice that all models struggle with reasoning classification tasks (MaRVL, XVNLI) where they obtain scores close to random guessing.
Table 6 additionally shows that Multilingual Open Flamingo trained on mOSCAR obtains much better results than the same model trained on WIT for equivalent training data seen during training [^22] We select the checkpoint of multilingual Open Flamingo trained on 35M documents and 70M captions to have fair comparison. (except for Multi30K benchmark) which means mOSCAR is better suited than WIT for training multilingual mLLMs. Eventually, Table 7 shows that we don’t face a drop in performances in English performances when training the model on 43 languages (multilingual Open Flamingo) in comparison to training it on the English subset of mOSCAR and English text-image pairs.
Additional comparison results with InternVL2 (Chen et al., 2024), Llava-NeXT (Li et al., 2024), PaliGemma (Beyer* et al., 2024) and Idefics2 (Laurençon et al., 2024) can be found in Appendix A.8.
@@ 5.2 Results
6 Conclusion
We introduced mOSCAR, a large-scale multilingual and multimodal dataset covering 163 languages and composed of 303M documents, 200B tokens and 1.15B images. We showed that mOSCAR is of good quality, diverse and could be used to train a multilingual and multimodal LLM. We also proved that training on mOSCAR led to the emergence of in-context learning capabilities in several languages. We eventually ensured that mOSCAR is as safe as possible by applying a series of filtering steps to remove NSFW and toxic content.
Limitations
We did not conduct any analysis of the biases of mOSCAR as this is challenging in a multilingual setting. As it is crawled from the Internet, it is indeed possible that mOSCAR reflects biases widespread on it. Training a model on mOSCAR must therefore be combined with additional alignment training steps to mitigate potential biases towards groups of people. Nevertheless, by its multilingual nature, mOSCAR is a step towards the inclusion of more languages, cultures, and people in accessing mLLMs.
@@ Acknowledgements
This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011014232R1, 2023-AD011014232 and 2023-AD011012254 made by GENCI. It was also partly funded by the last three authors’ chairs in the PRAIRIE institute funded by the French national agency ANR as part of the “Investissements d’avenir” programme under the reference ANR-19- P3IA-0001. We deeply thanks the Jean-Zay support team. We also thank Filip Šedivý for insightful discussions regarding the removal of CSAM, Thorn for having provided access to their CSAM detector, Zeeshan Khan for discussions regarding the training of the models and Victoria Le Fourner for having manually checked subsamples of NSFW images.
@@ Appendix A Appendix
Table LABEL:langs_stats shows the number of documents, the number of images, and the number of NLLB tokens for each language of mOSCAR.
A.2 Topic modeling
We analyzed the topic found in different language subsets of mOSCAR with BERTopic (Grootendorst, 2022). We first preprocessed $\sim$30K randomly sampled documents in tokens using stanza tokenizers (Qi et al., 2020), we then represent each document with their document embedding obtained with a multilingual sentence-bert232323sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 (Reimers and Gurevych, 2019) before running BERTopic to cluster the documents. Tables 10, 11, 12, 13 and 14 show some of the most salient English-translated strings representing each cluster and their weight (in percentage) in the subsets of mOSCAR. While many clusters are shared between subsets, we can observe that some of them are culturally-specific (i.e. they are related to the culture of the language subset and only appears on this subset). For instance, Table 14 shows topics of the Hindi subset and we observe a cluster related to Hindu gods described by the words ‘ganesha’, ‘shiva’ or ‘shani dev’. We also observe a cluster related to festivals hosted in India, it is described by the words ‘festival’, ‘dussehra’ or ‘janmashtami’. We can also observe that among clusters shared between several language subsets, some contain knowledge culturally-specific as exemplified by the cluster related to jewelry in the Hindi subset which contain the word ‘mangalsutra’ which is a jewel worn by men during weddings in India. Eventually, we can observe that the English subset contains a strong cluster related to the Christian religion. While a similar cluster appears in the Tosk Albanian subset, it is related to Islam. This shows the importance of a multilingual corpora to have data representing more cultures around the world.
A.3 Examples of documents
Figures 5, 7, 7, 8, 9 and 10 provide examples of documents found in mOSCAR for different languages. We can observe the diversity in terms of content of the documents. Figure 5 is a French document describing a specific shot of golf illustrated by images which provide very detailed description of images. Figure 7 shows a Russian document describing an Instagram account with pictures of food with animal heads which is reminiscent of a famous problem of computer vision consisting in classifying images between chihuahuas and muffins.